OpenStudy (anonymous):

If 4x-3y=12 represent line l and 12x+5y=0 represents line l'. find the equation of the locus of P under the following condition : 1) P is twice as far from l as from l' 2) is equidistant from l and l'

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what have you tried?

OpenStudy (anonymous):

P(x,y)

OpenStudy (amistre64):

2) should just be an average of the slopes, anchored to the point they have in common 1) is more interesting id say

OpenStudy (amistre64):

do you know any formulas that helps us define a points distance from a line?

OpenStudy (anonymous):

would that be the perpendicular one?\[|\frac{ Ax+By+C }{ \pm \sqrt{A^2+B^2} }|\]

OpenStudy (amistre64):

yes, as long as you are allowed to use it

OpenStudy (anonymous):

yea we shd be

OpenStudy (amistre64):

then set one distance to twice the other ... and simplify

OpenStudy (anonymous):

do we equate them? which one shd be twice though?

OpenStudy (amistre64):

l is 2 times longer than l' ... as given in the information

OpenStudy (amistre64):

l = 2l'

OpenStudy (amistre64):

and i believe your formula is for a line of the form: Ax + By + C = 0

OpenStudy (anonymous):

but... l is 2 times longer than l' ...i was thinking 2l=l'..lol

OpenStudy (amistre64):

d1 = distance to l d2 = distance to l' P is twice as far from l (d1) as from l' (d2) d1 = d2 + d2

OpenStudy (amistre64):